# Defect recombination origin of low energy excess in semiconductor detectors

### v1

@@ Abstract
When the detection threshold in semiconductor detectors is pushed to increasingly low energies, an “excess” signal of apparent energy release events below a few hundred eV is observed in several different kinds of detectors, hindering their sensitivity for rare event signals in this energy range. Using atomistic simulations with classical thermostat and quantum thermal bath, we show that this kind of signal is consistent with energy release from long-term recombination events of complex defects that can be formed by any kind of nuclear recoil radiation events. The recombination events are shown to have a very similar exponential dependence of energy as that observed in experiments. By detailed analysis of recombination events, we show that crossing a low energy barrier ($\lesssim$ 0.1 eV) can trigger larger energy releases in an avalanche-like effect. This explains why large energy release events can occur even down to cryogenic temperatures.
preprint: arxiv submission version, Aug 14, 2024
@@ I Introduction
One possible approach to detect light particle dark matter is by looking for low-energy recoil events in materials that cannot be attributed to conventional matter radiation events [1, 2, 3]. Similar methods are also used in coherent neutrino nucleus scattering experiments [4, 5]. This has motivated development of new detector technologies with lower and lower detection threshold.
@@
When the detection threshold in semiconductor detectors has been pushed to increasingly low energies, an “excess” signal of apparent energy release events below a few hundred eV is observed in several different kinds of detectors [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. This excess will mask any potential dark matter signal, and therefore understanding its origin and possible ways to mitigate it are questions of utmost importance. Several different explanations have been offered to this excess, for instance electronic noise [18], stress effects/crack relaxation [19, 20] or solid state effects [21]. However, some recent experiments disfavor an explanation due to external stress or radiation [22]. It seems plausible that there are several components and several different causes of the excess events.
Although there are thus many uncertainties regarding the origin of the effect, on a general level we note that the excess signal is observed in detectors consisting of different materials, manufacturing techniques and different electronics with a very similar exponential energy dependence $\propto\exp(-\alpha E)$, where $\alpha>0$ is aconstant. This suggests that there may be an inherent mechanism for the observed excess signal (consistent with the evidence for a solid state effects [21]).
@@
In this Article, we analyze the possibility that the excess signal could originate from recombination of nanometer-size range defect clusters. It is well established from both experiments [23, 24, 25, 26] and simulations [27, 28, 29, 30, 31] that keV recoils in semiconductor materials such as Si produce small disordered zones in the material. Even if the material is not irradiated explicitly, cosmic rays or decay of radioactive impurities that are always present at some concentration in a material and its immediate surrounding environment will produce similar damage pockets. Even if the initial kinetic energy of a radioactive decay event may be in the MeV range, the final damage will be produced by $\sim$ keV recoils, since collision cascades at much higher energies split up into subcascades [32, 28, 33].
@@
The nanometric damage regimes produced in such events are higher in potential energy than the pristine crystal structure, and hence are expected to thermally recombine towards the ground state at a time scale depending on temperature. Any recombination event will thus be associated with an energy release. While there have been previous works analyzing the recombination of radiation-induced disordered regions [27, 29, 34, 35], these have not systematically analyzed the energy distribution of the potential energy release events at cryogenic temperatures. In this article, we analyze this energy release systematically, and show that it leads to recombination events with an exponential energy dependence, very similar to that found in experiments.
@@ II Methods
To examine the production of disordered damage zones and their possible energy release on extended time scales, we used a five-stage simulation and analysis procedure: (i) simulation of typical radiation damage creation, (ii) simulation of thermal annealing of this damage on long time scales, (iii) quenching of configurations to 0 K to obtain potential energy release without effects from kinetic energy fluctuations, (iv) obtaining the energy release in recombination events from the time dependence of the potential energy in stage (iii), and (v) analysis of the time scale of annealing. These results are presented in section III. Since at cryogenic temperatures quantum mechanical zero point vibrations can affect atom dynamics, we also employed a quantum thermal bath (QTB) [36, 37] to assess how quantum effects affect the annealing. Results of these simulations are included along with the fully classical ones in the results, and discussed in section II.2. Finally, to understand why annealing is observed even at cryogenic temperatures, we also analyzed what kind of atomic movements in individual recombination events lead to energy release (section IV).
@@ II.1 Simulation and analysis stages
In all simulation stages (i-v), we used classical molecular dynamics [38]. Simulations were carried out mainly with the code PARCAS [39]. This code has been extensively used to examine radiation effects in a wide range of materials, and the results have been shown to agree well with experiments in many different measurable quantities [28, 40, 41, 42, 43, 44]. Some classical cases were also modelled with the widely used LAMMPS molecular dynamics code [45] and gave the same behaviour as the PARCAS results. The simulations with quantum thermal bath were all done with the LAMMPS codes; details on the specific quantum aspects of the modelling are given in subsection II.2. The analysis of annealing mechanisms in section IV were done with the LAMMPS code since this code has a built-in nudged elastic band (NEB) [46] feature that could be used in analyzing the energy barriers.
@@
The interatomic interactions were described with the Stillinger-Weber potential [47], that has been found to describe well a wide range of properties in Si when compared to experiments [48, 49], DFT calculations [31] and modern machine-learned potentials [49]. We also tested the Tersoff III potential [50], which gave very similar results regarding the energy release. In the following parts of the paper, unless otherwise mentioned the results shown are obtained with the Stillinger-Weber potential.
@@
In stage (i), we simulated the production of damage by recoil events in Si by simulating 5 keV cascades induced by self-recoils in bulk silicon. The energy of 5 keV is above the energy at which cascades break into subcascade in Si, i.e. the specific damage pockets produced by higher energy recoils are similar to those produced by 5 keV ones [28, 51]. The cascades were modelled following standard approaches [28, 52]. Briefly, the cascades were initiated in random directions in 3 dimensions with equal weighting of different directions by solid angle and the development of the cascade was simulated until the system cooled down back to the ambient temperature (0 K in the cascade phase). No temperature or pressure scaling was used in the collisional regions, and heat was removed at the system boundaries. Since in the current work, the emphasis is energy release from damage rather than the production process itself, we did not include the electronic stopping power [53] in the simulations. The simulation cell consisted of a cube with 16 unit cells in each dimension (i.e, a total of 32768 atoms) and periodic boundaries were used in all dimensions. This cell size was slightly smaller than what was normally used in cascade simulation, due to the necessity to be able to simulate annealing over long time scales. To ensure that no thermally induced annealing occurs in these damage generation events, the ambient temperature of the system was 0 K.
@@
In stage (ii), the damage produced in 5 keV cascades was annealed in thermal equilibrium runs over long time scales. In most cases we used 3 ns as the annealing time, but we tested annealing time up to 30 ns. The annealing runs were done for temperatures in the range of 0.05 - 800 K. To enable analysis of damage recombination events that can occur quite rapidly, the atom coordinates of all atoms were output every 2 ps throughout the entire simulation. In the annealing runs, Berendsen temperature control was used on all atoms with a time constant $\tau_{T}=300$ fs and pressure control to 0 pressure with a time constant of $\tau_{P}=3000$ fs.
@@
In stage (iii), the atom coordinates output from the annealing run were extracted every 2 ps, each one separately, as input coordinates for quenching. The quenching simulations were performed efficiently using Berendsen temperature control [54] towards 0 K while zeroing the velocity of any atom $i$ on which the force vector $\vec{F}_{i}$ was in opposite direction to the velocity vector $\vec{v}_{i}$, i.e. if $\vec{F}_{i}\cdot\vec{v}_{i}<0$. Using this efficient approach, a quenching time of 1 ps was enough to get the system temperature down below 0.01 K. The quenching was done every time when the atom coordinates were output in the 3 ns (or longer) annealing runs, and the final potential energy was recorded after quenching.
@@
In stage (iv), the potential energy vs. time in stage (iii) was analyzed with respect to energy release. After quenching to 0 K, the energy release events were found to be clearly distinct from minor random fluctuations in potential energy that was $<0.5$ eV. To get the total energy release from one recombination event, when energy release $>0.5$ eV occurred on subsequent 2 ps time intervals, the energy was summed up as the total release in the same event. Subsequently a statistics of the magnitude of energy release was made. This can be compared with the experimental observations of energy release.
In stage (v), the annealing time scale was analyzed by fitting a stochastic time scale law to the data on the time dependence of annealing obtained from molecular dynamics. This allowed us to quantitatively show that annealing of amorphous pockets can occur even at extremely low temperatures relevant for superconducting transition detectors.
@@
For defect analysis, the number of defects in the simulation cells was quantified with the Wigner-Seitz (WS) method [28], which is a geometric method completely independent of the potential energy analysis. This measure is used to quantify the number of defects because its definition has no adjustable parameters, and it gives defect numbers roughly consistent with the simple Kinchin-Pease damage prediction in Si [28, 52]. However, considering the complex nature of the damage pockets in Fig. 2, it is clear that in Si “the number of WS defects” does not correspond to any geometrically simple point defect type such as an isolated vacancy or interstitial. Hence we use the term “the number of WS defect pairs” as the notation of this defect measure.

### v2

# Spontaneous damage annealing reactions as a possible source of low energy excess in semiconductor detectors

@@ Abstract
In semiconductor detectors designed for capturing dark matter particles or neutrinos, when the detection threshold is constantly improved to increasingly low energies, an “excess” signal of apparent energy release events below a few hundred eV is observed in several different kinds of detectors. This becomes a big obstacle to the observation of actual dark matter signals, hindering the detectors’ sensitivity for rare events in this energy range. Using atomistic simulations with a classical thermostat and a quantum thermal bath, we show that this kind of signal is consistent with energy release from long-term annealing events of complex defects that can be formed by any kind of nuclear recoil radiation events. Such energy releases are shown to have a very similar exponential dependence on energy release magnitudes as that observed in experiments. By detailed analysis of the annealing events, we show that crossing very low energy barriers can trigger larger energy releases in an avalanche-like effect. This explains why large energy release events can occur even down to cryogenic temperatures, where the significant migration of point defects in silicon is hardly ever possible.
preprint: Under revision in Physical Review Materials (2025)
@@ I Introduction
One possible approach to detect light particle dark matter is by looking for low-energy recoil events in materials that cannot be attributed to conventional matter radiation events [1, 2, 3]. Similar methods are also used in coherent neutrino nucleus scattering experiments [4, 5]. This strong demand from physics research has motivated the consistent development of new detector technologies with lower and lower detection threshold.
@@
When the detection threshold in semiconductor detectors has been pushed to increasingly low energies, an “excess” signal of apparent energy release events below a few hundred eV is observed in several different kinds of detectors [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], being able to mimic or cover a possible DM signal. Sharing a similar energy range with possible dark matter recoil signals, this excess will mask any potential dark matter signal, and therefore understanding its origin and finding out possible ways to mitigate or even eliminate it are questions of utmost importance. Several different explanations have been offered to this excess, for instance electronic noise [18], stress effects/crack relaxation [19, 20] or solid state effects [21]. However, some recent experiments disfavor an explanation due to intrinsic effects in the crystal or external radiation [17]. It seems somewhat convincing that there are several components and several different causes of the excess events.
Although there are thus many uncertainties regarding the origin of the effect, on a general level we note that the excess signal is observed in detectors consisting of different materials, manufacturing techniques and different electronics with a very similar exponential energy dependence $\propto\exp(-\alpha E)$, where $\alpha>0$ is a constant. This outward phenomenon suggests that there may be an inherent mechanism for the observed excess signal, due to this very obvious similarity in the thermodynamic patterns of the energy release spectra in distinct experiments. Hypothesizing the existence of a common origin is also consistent with the evidence pointing to a solid-state effect [21] as the source of the signal. This inherent mechanism should not depend on any specific choices in detector materials, the specific fabrication techniques of detector materials, or the design of the electric circuits in the detectors, and cannot be weakened by the modifications in these very aspects in the detectors. This character of invariance really attracts researchers to dig deeper for the identification of the actual source behind the scene. Even if the investigations are still ongoing and have not reached any consensus on this, the academic gains along this discovery path have already been quite fruitful [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], having very profound influence in dark matter detection experiments and science or engineering technology as a whole.
@@
In this article, we analyze the possibility that the excess signal could originate from local annealing of nanometer-size range defect clusters. It is well established from both experiments [22, 23, 24, 25] and simulations [26, 27, 28, 29, 30] that keV recoils in semiconductor materials such as Si produce small disordered zones in the material. Even if the material is not irradiated explicitly, cosmic rays or decay of radioactive impurities that are always present at some concentration in a material and its immediate surrounding environment will produce similar damage pockets. Even if the initial kinetic energy of a radioactive decay event may be in the MeV range, the final damage will be produced by $\sim$ keV recoils, since collision cascades at much higher energies split up into sub-cascades [31, 27, 32].
@@
The nanometric damage regimes produced in such events are higher in potential energy than the pristine crystal structure, because of the existence of overcoordination of atoms, undercoordination of atoms and all kinds of deviations from the tetrahedral structures in the disordered damaged region. Hence, these damages are expected to thermally anneal towards the ground state at a time scale depending on temperature. The annealing, which may occur stepwise, will thus be associated with energy release events. We note that in the traditional theory of radiation damage, the damage is assumed to be distributed in well-defined isolated interstitials and vacancies, which have a well-defined formation energy and hence if they meet and recombine, the energy release would be always well-defined. However, numerous studies show that the damage in covalently bonded semiconductors is much more complex [26, 27, 33] and hence there may be a wide range of different energy release values. For this reason, we tend to avoid using the word “recombination” to describe the energy release events. While there have been previous works analyzing the annealing of radiation-induced disordered regions [26, 28, 34, 35, 36, 37], these have not systematically analyzed the energy magnitude distribution of the potential energy release events in annealing. Also, very, very few of them, if any, looked into what happened at cryogenic temperatures, as this is not a relevant temperature regime in industrial applications like semiconductor chip manufacturing. In this article, we analyze this energy release magnitudes systematically, and show that it leads to energy release events with an exponential energy dependence, very similar to that found in experiments.
@@ II Methods
To examine the production of disordered damage zones and their possible energy release on extended time in annealing, we used a five-stage simulation and analysis procedure: (i) simulation of typical radiation damage creation, (ii) simulation of thermal annealing of this damage on long time scales, (iii) quenching of configurations to 0 K to obtain potential energy release without effects from kinetic energy fluctuation, (iv) obtaining the energy release in annealing events from the time dependence of the potential energy in stage (iii), and (v) analysis of the time scale of annealing. These results are presented in section III. Since at cryogenic temperatures quantum mechanical zero point vibrations can affect atom dynamics, we also employed the quantum thermal bath (QTB) [38, 39] to assess how quantum effects affect the annealing. The basic ideas of utilizing QTB are introduced in section II.2, and the results of these simulations are included along with the fully classical ones in the following discussions. Finally, to understand why annealing is observed even at cryogenic temperatures, we also analyzed what kind of atomic movements in individual annealing events (energy release events) actually lead to the energy release (section IV and section V).
@@ II.1 Simulation and analysis stages
In all simulation stages (i-v), we used classical molecular dynamics [40]. As for classical molecular dynamics simulations, several of them were carried out with the code PARCAS [41]. This code has been extensively used to examine radiation effects in a wide range of materials, and the results have been shown to agree well with experiments in many different measurable quantities [27, 42, 43, 44, 45, 46]. In comparison to other popular molecular dynamics simulators (for example, LAMMPS), PARCAS has a unique advantage of possessing quite a number of choices for accurate repulsive interatomic potentials, which are important when doing careful research on irradiation damage creation by irradiation cascades. Besides the classical simulations conducted by PARCAS, a number of other classical cases were modeled with the widely used LAMMPS molecular dynamics code [47] and gave the same behaviour as the PARCAS results. The simulations with the quantum thermal bath (QTB) were all done with the LAMMPS codes, since PARCAS does not include this QTB feature. Details on the specific quantum aspects of the modelling are given in subsection II.2. The analysis of annealing mechanisms in section IV were done with the LAMMPS code since this code has a built-in climbing image nudged elastic band (CI-NEB) [48] feature that could be used in analyzing the energy barriers, which is an essential part in the evaluation of kinetics of annealing events.
@@
The interatomic interactions between silicon atoms were described with the Stillinger-Weber potential (SW potential) [49], that had been found to describe well a wide range of properties in Si when compared to experiments [50, 51], DFT calculations [30] and results from modern machine-learned potentials [51]. We also tested the Tersoff III potential [52], which gave very similar results regarding the energy release behavior. The Tersoff III potential was used as a supplement in our research due to a few reasons. First, the development of this and the SW potential potential were two completely independent processes, so Tersoff results can serve as a comparison with SW results. Second, Tersoff III potential has been used quite intensively in irradiation cascade simulations for silicon [27]. Moreover, in the PARCAS code, this potential has been augmented by a DMol repulsive potential [53, 54] using the Fermi function joining formalism [55].
In the cascade runs done by PARCAS code, the Stillinger-Weber potential was also augmented by the DMol repulsive potential [53, 54]. The cascade runs carried out with the LAMMPS code did not use the repulsive potential, since the code does not support adding a repulsive part to the Stillinger-Weber potential. However, we ran some of the LAMMPS annealing runs based on the PARCAS cascade results, using the output geometries from PARCAS cascade siulations as the input for annealing simulations conducted by LAMMPS. The results on the energy release behavior were in nature identical with pure LAMMPS runs within the statistical error bars. In the following parts of the paper, unless otherwise mentioned, the results shown are obtained with the Stillinger-Weber potential, and it is our main tool in this investigation.
@@
In stage (i), we simulated the production of damage by recoil events in Si by simulating 5 keV cascades induced by self-recoils in bulk silicon. The energy of 5 keV is above the energy at which cascades break into sub-cascade in Si, i.e. the specific damage pockets produced by higher energy recoils are similar to those produced by 5 keV recoils [27, 56]. The cascades were modelled following standard approaches [27, 57]. Briefly, the cascades were initiated in random directions in 3 dimensions (polar angle $\theta$ and azimuthal angle $\phi$) with equal weighting of different directions in terms of the solid angle and the development of the cascade was simulated until the system was cooled down back to the ambient temperature, namely, 0 K in the cascade phase. No temperature or pressure scaling was used in the collisional regions, and heat was removed at the system boundaries, using a Berendsen thermostat applied to a layer of atoms on the boundaries. Importantly, to properly describe high-energy interactions, we merged the DFT-based “DMol” potential at short interatomic separations [53, 54] using the Fermi function smooth joining formalism [58]. For different temperature levels in annealing (described in the following parts), the number of recoil cases considered ranged between 10 and 1000, with the highest statistics being used for the lowest-temperature simulations. In these runs of lower temperatures, the annealing events were rarer, so a higher statistics was needed. The number of cases simulated is reflected in the statistical error bars.
In the current work, the emphasis is the energy release from damage annealing processes, and the damage production process itself is not our focus. Therefore, we did not include the electronic stopping power [59] in the simulations. The simulation cell was a cube of silicon atoms with 16 unit cells in each dimension (i.e, a total of $16\times 16\times 16\times 8=32768$ atoms) and periodic boundaries were used in all dimensions. This cell size was slightly smaller than what had been normally used in cascade simulations [27], due to the necessity to be able to simulate annealing over long time scales at a reasonable computational cost. Although the simulation cell is smaller than the ones used usually in cascade simulations, it is very unlikely for the cascade damage region to touch itself via the periodic boundaries, since in Si the cascade region is broken into many sub-cascade regions when the recoil energy is 5 keV, and the recoil energy can be distributed in the newly formed sub-cascades. To avoid the occurrence of thermally induced defect annealing in cascade simulations, or at least to a great extent alleviate them, we set the ambient temperature of the system to zero by using a Berendsen thermostat on boundary regions to remove heat from the boundaries. The target temperature of this boundary thermostat was 0 K and the temperature damping parameter was $\tau_{T}=0.1ps$.
In stage (ii), the damage produced in 5 keV cascades was annealed in thermal equilibrium runs over long time scales. Annealing simulations were separated into two big groups. One group utilized PARCAS as the simulation platform, and consequently, this group of annealing simulations only used Berendsen thermostat, since PARCAS only had Berendsen method as the thermostat. All PARCAS annealing simulations in this group lasted for 3 ns. The other group used LAMMPS as the simulation platform. In this group of annealing simulations, some runs were conducted via Berendsen thermostat, while others were carried out using the quantum thermal bath (QTB thermostat). When using LAMMPS, no matter which thermostat we used, QTB or Berendsen, the annealing runs were conducted for 10 ns. The annealing runs were done for temperatures in the range of 0.05-800 K. No matter which simulator we used, PARCAS or LAMMPS, in these annealing runs, the thermostat is always applied to all atoms in the system.
In the annealing runs, a thermostat was applied to all the atoms. When Berendsen temperature control was used, the temperature damping parameter was $\tau_{T}=300fs$. When using the QTB thermostat, the temperature damping parameter was the same. Besides the thermostat, a Berendsen barostat was also applied to the system for pressure control, regardless of the type of the thermostat. The pressure damping parameter of the barostat was $\tau_{P}=3000fs$.
The number of annealing runs performed varied, from 7 for the 500 K case to 60 for the 200 K case in the classical simulations, and several hundreds for the quantum thermal bath ones. The highest statistics was used for the lowest temperature cases, since at these temperatures the annealing events were rarer, requiring a higher number of statistics. Another reason for this variation was that the higher temperatures were not of particular interest, and hence a smaller statistics was sufficient for them. The statistics settings in the classical simulations and the quantum thermal bath simulations, namely, the number of cases at each temperature level and each thermostat choice, were reflected in the error bars. The number of recoil cases at each temperature level, interatomic potential, and choice of thermostat, is given in Table 1.
@@
To enable the analysis in the very next simulation stage on damage annealing events that could occur quite rapidly, the atom coordinates of all atoms were output every 2 ps throughout the entire simulation.
Table 1: The number of recoil cases used in cascade annealing simulations at different temperatures and thermostats. QTB stands for the quantum thermal bath, and Berendsen stands for the classical Berendsen thermostat. SW stand for Stillinger-Weber interatomic potential, and Tersoff stands for Tersoff III interatomic potential.
Temperature (K) Thermostat Number of Cases Potential
0.05 QTB 800 SW
10 QTB 600 SW
25 QTB 600 SW
50 QTB 400 SW
100 QTB 200 SW
100 Berendsen 200 SW
125 Berendsen 200 SW
150 QTB 200 SW
150 Berendsen 200 SW
200 QTB 200 SW
200 Berendsen 200 SW
300 QTB 200 SW
300 Berendsen 200 SW
300 Berendsen 40 Tersoff
@@ 50 QTB 400 SW
In stage (iii), the atom coordinates output from the annealing simulations were extracted every 2 ps, each one separately, as input coordinates for quenching. The quenching simulations were performed efficiently using Berendsen temperature control [60] towards 0 K while zeroing the velocity of any atom $i$ on which the force vector $\vec{F}_{i}$ was in opposite direction to the velocity vector $\vec{v}_{i}$, i.e. if $\vec{F}_{i}\cdot\vec{v}_{i}<0$. This quenching method is called “quickmin” or “damped dynamics method” in the manual of LAMMPS [61]. Using this efficient approach, a quenching time of 1 ps was enough to get the system temperature down below 0.01 K. The quenching was done every time when the atom coordinates were output in the 3 ns annealing runs with Berendsen thermostat, and in the 10 ns annealing runs with QTB thermostats, that is to say, every 2 ps throughout the simulations. The final potential energy yielded by this energy minimization process was recorded after quenching.
@@
In stage (iv), the potential energy vs. time in stage (iii) was analyzed with respect to energy release. After being quenched to 0 K, the energy release events were found to be clearly distinct from random thermal fluctuations in potential energy, that were <0.5<0.5 eV. Moreover, when comparing the energy release with the time dependence of the number of Wigner-Seitz defects in the system, one can find out that such energy release could be related to changes in atomic structure. To get the total energy release from one annealing event, when energy release >0.5>0.5 eV occurred on subsequent 2 ps time intervals, the energy release values were summed up as the total release in the same event. Subsequently, a statistics of the magnitude of energy release was made. This can be compared with the experimental observations of energy release, which is the low energy excess signal.
In stage (v), the annealing time scale was analyzed by fitting a stochastic time scale law to the data on the time dependence of annealing obtained from molecular dynamics. To do this fitting, the potential energy curves of all the simulations at each temperature level, from each particular choice of thermostat, Berendsen or QTB, were summed up and averaged, yielding an “average” potential energy curve at this temperature level. Then, the part of the average curve from 1 ns to the end of the annealing simulations (3 ns for PARCAS simulations and 10 ns for LAMMPS simulations) was selected, and this latter, relatively long-term part, was fit by an exponential function. The data between 0 ns to 1 ns was not used in the time constant fitting, since the most unstable defects are likely to be annealed first, causing an unrealistic amount of energy release. This allowed us to quantitatively show that annealing of amorphous pockets can occur on long time scales, even at extremely low temperatures relevant for superconducting transition detectors.
@@
For defect analysis, the number of defects in the simulation cells was quantified with the Wigner-Seitz (WS) method [27], a geometric method completely independent of the potential energy analysis. This measure is used to quantify the number of defects because its definition has no adjustable parameters, and it gives defect numbers roughly consistent with the simple Kinchin-Pease damage prediction in Si [27, 57]. However, considering the complex nature of the damage pockets in Figure 1, it is clear that in Si, “the number of WS defects” does not correspond to any geometrically simple point defect type such as an isolated vacancy or interstitial. Hence we use the term “the number of WS defect pairs” as the notation of this defect measure. Also, when analyzing the microscopic mechamisms of energy release events, the DXA analysis in Ovito [62, 63] was also used to examine the size evolution of the amorphous pockets in silicon, This approach is a quantitative and useful approach for the investigation of the crystallization of amorphous pockets under thermal activation. To use DXA analysis on damaged silicon systems in this project, we utilized the Python package of Ovito, and made programs to calculate the number of atoms in the amorphous phase in the annealing processes. Besides being as handy as WS defect analysis, there is another reason why we used DXA despite the existence of WS defect analysis in our research. In Ovito, DXA analysis can generate a “defect mesh” to visually mark the boundaries of amorphous pockets in silicon, being very useful when finding out the precise locations of energy release events (in section V).
Figure 1: Cross-sectional view of the time evolution of damage formation in a 5 keV cascade in Si, as modeled with the Stillinger-Weber potential and a classical thermostat. In these simulations, one atom in the cell was given a recoil energy of 5 keV towards the center of the cell, leading to a collision cascade producing damage. For clarity, only atoms in a 2 unit cell thick cross section in the center of the cell are shown. Atoms are of course colliding and damage is produced also in other atom layers than those shown. The top part shows the time evolution of the cascade at the specific times 0, 0.1, 0.4 and 1 ps, and the larger frame at the bottom the final damage state at 10 ps. This cascade was ran at 0 K ambient temperature to ensure there is no thermal annealing. The precise size of the plotted simulation cell region is 81.2 Å in the $x$ (horizontal), 71.3 Å in the $y$ (vertical) and 10.86 Å in the $z$ (out of plane) direction. The colors indicate the potential energy $E_{\rm pot}$ of the atoms relative to the ground state energy $E_{\rm pot,0}$, as indicated in the legend on the top right.